How do you Divide mixed fractions? So 5 1/4 divided by 4 3/8? Would the answer be 9 2/3?
Answer by nerdybill(7384)

How do you Divide mixed fractions? So 5 1/4 divided by 4 3/8?
.
First, convert both to an improper fraction.
.
5 1/4 = 5 + 1/4 = 20/4 + 1/4 = 21/4
and
4 3/8 = 4 + 3/8 = 32/8 + 3/8 = 35/8
.
so 5 1/4 divided by 4 3/8 can be written as:
(21/4)/(35/8)
multiplying numerator and denominator by 8 we get:
42/35
or
1 + 7/35
or
1 + 1/5
or
1 1/5 (answer)